OpenStudy (anonymous):

How many solutions does this system of equations have? 3x – 4y = 12 y = 4x – 3 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

you make no sense

OpenStudy (anonymous):

you can use substitution.

OpenStudy (anonymous):

How many solutions does this system of equations have though not solving the equation

OpenStudy (anonymous):

solving the equation is one way to find out

OpenStudy (anonymous):

oh ok sorry i'm still asleep

OpenStudy (anonymous):

the first equation can be re-arranged 3x – 4y = 12 3x = 12 + 4y 4y = 3x - 12 y = 3/4 -12

OpenStudy (anonymous):

so we have these two equations: y = 3/4x -12 y = 4x – 3

OpenStudy (anonymous):

you can't see at one sight how many solutions they have

OpenStudy (anonymous):

so I can just solve the system instead :) 3x – 4y = 12 ,plug y=4x-3 3x - 4(4x-3) = 12 3x -16x +12 = 12 3x-16x = 0 -13 x = 0 x = 0.....

OpenStudy (anonymous):

y = 4x – 3 y = -3 3x – 4y = 12 -4y = 12 y = -3 there seems to be a solution

OpenStudy (anonymous):

Oh I see haha thank you

OpenStudy (anonymous):

no problem, I'm still trying if elimination gives more information 3x – 4y = 12 y = 4x – 3 – 4y = -3x + 12 y = 4x – 3 – 4y = -3x + 12 4y = 16x – 12 ------------ 0 = -9x

OpenStudy (anonymous):

looks like x really is 0........and y=-3

OpenStudy (anonymous):

ok i'm once again confuzzled

OpenStudy (anonymous):

its because we dont follow rules........... do you know any rules for the outcome of substitution? I believe ifyou substitute two equations and you get a distinct value, then the system has 1 solution

OpenStudy (anonymous):

at least if the solution works in both equations, which we have checked :)

OpenStudy (anonymous):

y - x = 9 2 y - 2x = 18 this system has infinite solutions what happens when subsituting? the first equation can be written as y = 9 + x, then substitute: 2 ( 9+x ) -2x = 18 18 + 2 x - 2 x = 18 18 = 18 look at this: if you get a solution like this, then it has infinite solutions

OpenStudy (anonymous):

oh ok thank you but i'm litterally falling asleep I'll just do this stuff after a couple hours thank you so much